What should be the tension in a wire so that the speed of the transverse wave travelling through it equals the speed of sound?

Grade:12

1 Answers

Riddhish Bhalodia
8 years ago
we know that speed of a transverse wave relates to the force with the following relation
v = \sqrt{\frac{F}{\mu}} \quad where \quad \mu = M/L
Force here is the tension in the wire and soundwaves travel at 340m/s = v
so you can solve it